OpenStudy (anonymous):

The y-intercept of the line whose equation is 3x + 2y = 7 -3/2 correct? 7/3 7/2

OpenStudy (anonymous):

no that is not correct, try making x=0

OpenStudy (anonymous):

3x + 2y = 7 In slope-intercept form: 2y = -3x + 7 y = (-3/2)x + 7/2 y-intercept = 7/2

OpenStudy (anonymous):

ok. Thanks